# Identifying Policy Gradient Subspaces

### v1

@@ 4.3 The high-curvature subspace changes slowly throughout the training
So far, we have verified that the gradients of the actor and critic losses optimized by PPO and SAC lie to a large extent in the subspace spanned by the top eigenvectors of the Hessian with respect to the current parameters. However, even though there are relatively efficient methods for computing the top Hessian eigenvectors without explicitly constructing the Hessian matrix, calculating these eigenvectors at every step would be computationally expensive. Ideally, we would like to identify a subspace once that remains constant throughout the training. In practice, the gradient subspace will not stay exactly the same during the training, but if it changes slowly, it is possible to reuse knowledge from earlier timesteps and update the subspace at a lower frequency. To that end, we investigate condition iii) by calculating the subspace overlap, defined by Gur-Ari et al. (2018). The subspace overlap between subspaces identified at different timesteps and is specified as
@@
where is the th largest eigenvector at timestep . denotes the projection matrix from the full parameter space to the high-curvature subspace, identified at timestep . Similar to Equation 5, the criterion measures how much of the original vector is preserved during the projection into the subspace. For the subspace overlap, however, we use the projection matrix at timestep not to project the gradient but rather project the Hessian eigenvectors that span the high-curvature subspace identified at a later timestep of the training. This criterion, thus, measures how much the gradient subspace changes between these timesteps. Note that we assume the eigenvectors to be normalized to one and therefore do not normalize by the eigenvector length.
@@
Gur-Ari et al. (2018) showed in the supervised setting that the gradient subspace stabilizes only after some initial update steps. Therefore, we choose the timestep at which we identify the projection matrix as since this is still relatively early in the training, but the gradient subspace should already have stabilized reasonably well. As in Section 4.2, we use as subspace dimensionality. The analysis results in Figure 3 show that the subspace overlap reduces the further apart the two timesteps and are, but in all cases, the subspace overlap remains considerably above zero, implying that information of previous subspaces can be reused at later timesteps. If the two timesteps are close to each other, the overlap is considerable. Similar to the gradient subspace fraction in Section 4.2, the subspace overlap is often more pronounced for the critic than the actor.
The SAC results for Ant and Walker2D show a small drop in the subspace overlap at 1.1 million steps. At this timestep, the data from timestep in the replay buffer is replaced completely by new data. The drop, therefore, is an indicator that a changing data distribution has a negative effect on the subspace overlap. Note that this is not captured by the plots for the Finger-spin and LunarLanderContinuous tasks since the algorithms converge before reaching 1 million timesteps, and, thus, we only plot the analysis results up to this timestep. Moreover, the subspace overlap tends to be larger for SAC than PPO for timesteps close to the original timestep but smaller for timesteps further away. We hypothesize that this is due to two opposing effects: 1) The data distribution changes more slowly, which we found to be beneficial in terms of subspace overlap. 2) The networks of SAC are significantly larger and, hence, there are more dimensions in which the subspace might change. The results above indicate that effect 1) has a stronger influence early, while effect 2) dominates later on.
@@ 5.2 Potential of gradient subspaces in reinforcement learning
While the network architectures commonly used in reinforcement learning are small compared to the models used in other fields of machine learning, the dimensionality of the optimization problem is still considerable. Widely used optimizers, like Adam (Kingma & Ba, 2014), often rely only on gradient information, as computing the Hessian at every timestep would be computationally very demanding in high dimensions. However, in Section 4.1, we have seen that the optimization problem is ill-conditioned. Second-order methods, like Newton’s method, are known to be well-suited for ill-conditioned problems (Nocedal & Wright, 1999). With the insights of this paper, it seems feasible to reduce the dimensionality of the optimization problems in RL algorithms by optimizing in the low-dimensional subspace instead of the original parameter space. The low dimensionality of the resulting optimization problems would enable computing and inverting the Hessian matrix efficiently and make second-order optimization methods feasible.
@@ 6 Reproducibility
We applied our analyses to well-known, publicly available benchmark tasks and utilized popular, publicly available implementations of the RL algorithms from Stable-Baselines3 (Raffin et al., 2021). Further experimental details like the learning curves of the algorithms and the fine-grained analysis results for the entire training are displayed in Appendices B and C, respectively. Furthermore, we published the code for the analyses to allow reproduction of our experiments here.
Appendix A Derivation of the gradient subspace fraction objective
In this section, we derive Equation 6. Recall that is the matrix of the largest Hessian eigenvectors, which is semi-orthogonal. We use to denote the projection of the gradient into the subspace and back to its original dimensionality. In the following derivation, we drop the subscript of matrix for ease of notation.
Appendix B Learning curves
Appendix C Detailed analysis results for all tasks
(a) Ant, actor
(b) Ant, critic
(c) Ball_in_cup, actor
(d) Ball_in_cup, critic
(e) Bip.Walker, actor
(f) Bip.Walker, critic
(g) FetchReach, actor
(h) FetchReach, critic
(i) Finger-spin, actor
(j) Finger-spin, critic
(k) HalfCheetah, actor
(l) HalfCheetah, critic
(m) Hopper, actor
(n) Hopper, critic
(o) LunarLander, actor
(p) LunarLander, critic
(q) Pendulum, actor
(r) Pendulum, critic
(s) Reacher, actor
(t) Reacher, critic
(u) Swimmer, actor
(v) Swimmer, critic
(w) Walker2D, actor
(x) Walker2D, critic
(a) Ant, actor
(b) Ant, critic
(c) Ball_in_cup, actor
(d) Ball_in_cup, critic
(e) Bip.Walker, actor
(f) Bip.Walker, critic
(g) FetchReach, actor
(h) FetchReach, critic
(i) Finger-spin, actor
(j) Finger-spin, critic
(k) HalfCheetah, actor
(l) HalfCheetah, critic
(m) Hopper, actor
(n) Hopper, critic
(o) LunarLander, actor
(p) LunarLander, critic
(q) Pendulum, actor
(r) Pendulum, critic
(s) Reacher, actor
(t) Reacher, critic
(u) Swimmer, actor
(v) Swimmer, critic
(w) Walker2D, actor
(x) Walker2D, critic
(a) Ant, actor
(b) Ant, critic
(c) Ball_in_cup, actor
(d) Ball_in_cup, critic
(e) Bip.Walker, actor
(f) Bip.Walker, critic
(g) FetchReach, actor
(h) FetchReach, critic
(i) Finger-spin, actor
(j) Finger-spin, critic
(k) HalfCheetah, actor
(l) HalfCheetah, critic
(m) Hopper, actor
(n) Hopper, critic
(o) LunarLander, actor
(p) LunarLander, critic
(q) Pendulum, actor
(r) Pendulum, critic
(s) Reacher, actor
(t) Reacher, critic
(u) Swimmer, actor
(v) Swimmer, critic
(w) Walker2D, actor
(x) Walker2D, critic
(a) Ant, actor
(b) Ant, critic
(c) Ball_in_cup, actor
(d) Ball_in_cup, critic
(e) Bip.Walker, actor
(f) Bip.Walker, critic
(g) FetchReach, actor
(h) FetchReach, critic
(i) Finger-spin, actor
(j) Finger-spin, critic
(k) HalfCheetah, actor
(l) HalfCheetah, critic
(m) Hopper, actor
(n) Hopper, critic
(o) LunarLander, actor
(p) LunarLander, critic
(q) Pendulum, actor
(r) Pendulum, critic
(s) Reacher, actor
(t) Reacher, critic
(u) Swimmer, actor
(v) Swimmer, critic
(w) Walker2D, actor
(x) Walker2D, critic
Appendix D Impact of suboptimal hyperparameters
Hyperparameters typically have a significant impact on the learning performance of policy gradient algorithms (Paul et al., 2019). In Section 4.2, we analyzed the gradient subspace for training runs with tuned hyperparameter configurations. However, which hyperparameters work well for a given problem is typically not known a priori. Finding good hyperparameters often involves running numerous RL trainings, which might be infeasible when training on real-world tasks. For our insights to be valuable for such real-world settings, it is crucial that suitable gradient subspaces can also be identified for training runs with suboptimal hyperparameters. Therefore, in this section, we investigate the robustness of the gradient subspace with respect to suboptimal hyperparameters. To get a set of hyperparameters that are realistic but potentially suboptimal, we sample hyperparameter configurations from the ranges that are found frequently in the tuned hyperparameters of RL Baselines3 Zoo (Raffin, 2020). Particularly, we draw the samples from the sets given in Table 1. Note that these are also the bounds we use when sampling configurations for hyperparameter tuning.
| Algorithm | Hyperparameter | Values | Logscale |
| PPO | learning_rate | yes | |
| batch_size | no | ||
| n_steps | no | ||
| n_epochs | no | ||
| gamma | no | ||
| gae_lambda | no | ||
| clip_range | no | ||
| ent_coef | yes | ||
| net_arch | no | ||
| SAC | learning_rate | yes | |
| batch_size | no | ||
| train_freq | no | ||
| gamma | no | ||
| tau | no | ||
| learning_starts | no | ||
| net_arch | no |
We sampled a total of 50 hyperparameter configurations per algorithm and task. Figure 9 displays the distribution over the maximum cumulative reward that these hyperparameter configurations achieve. As expected, the variance in the performance is large across the hyperparameter configurations. While some configurations reach performance levels comparable to the tuned configuration, most sampled configurations converge to suboptimal behaviors.
To display the values of the gradient subspace fraction and the subspace overlap compactly for all configurations, we compress the results for each training run to a single scalar. For the gradient subspace fraction, we compute the mean of the criterion over the timesteps. Regarding the subspace overlap, we choose the early timestep in accordance to the experiments in Section 4.3. However, taking the mean over the remaining timesteps would not be faithful to how the subspace would likely be utilized in downstream applications. Instead of identifying the gradient subspace once and using it for the rest of the training, such methods would likely update the gradient subspace after a given number of gradient steps. The rate of this update would likely depend on the exact way that the subspace is utilized. For this analysis, we choose , so that the timestep difference of steps is significantly shorter than the total length of the training, but the number of gradient steps is large enough that the subspace could change.
The results for the gradient subspace fraction and the subspace overlap for random hyperparameters are visualized in Figure 10 and Figure 11, respectively. Figure 10(a) shows for SAC’s actor that the gradient is well contained in the subspace for all random hyperparameter configurations. For PPO’s actor, the spread in the gradient subspace fraction is significantly higher. As shown in Figure 10(b), the gradient subspace fraction for the critic is very high for all hyperparameter configurations. These results suggest that gradient subspaces can be utilized in SAC independently of the hyperparameter setting, while PPO’s actor might require more considerations.
Figure 11 shows that there is a significant spread in the subspace overlap for the random hyperparameter configurations, which indicates that potential downstream applications might need to update the gradient subspace more frequently, depending on the hyperparameters.

### v3

@@ 4.3 The high-curvature subspace changes slowly throughout the training
So far, we have verified that the gradients of the actor and critic losses optimized by PPO and SAC lie to a large extent in the subspace spanned by the top eigenvectors of the Hessian with respect to the current parameters. However, even though there are relatively efficient methods for computing the top Hessian eigenvectors without explicitly constructing the Hessian matrix, calculating these eigenvectors at every step would be computationally expensive. Ideally, we would like to identify a subspace once that remains constant throughout the training. In practice, however, the gradient subspace will not stay exactly the same during the training, but if it changes slowly, it is possible to reuse knowledge from earlier timesteps and update the subspace at a lower frequency. To that end, we investigate condition iii) by calculating the subspace overlap, defined by Gur-Ari et al. (2018). The subspace overlap between timesteps and is defined as
@@
where is the th largest eigenvector at timestep . denotes the projection matrix from the full parameter space to the high-curvature subspace, identified at timestep . Similar to Equation 5, the criterion measures how much of the original vector is preserved during the projection into the subspace. For the subspace overlap, however, we use the projection matrix at timestep not to project the gradient but rather project the Hessian eigenvectors that span the high-curvature subspace identified at a later timestep of the training. This criterion, thus, measures how much the gradient subspace changes between these timesteps. Note that we assume the eigenvectors to be normalized to one and therefore do not normalize by their length.
@@
Gur-Ari et al. (2018) showed in the supervised setting that the gradient subspace stabilizes only after some initial update steps. Therefore, we choose the timestep at which we initially identify the subspace as since this is still relatively early in the training, but the gradient subspace should already have stabilized reasonably well. We evaluate the subspace overlap criterion every timesteps until timestep . This interval covers a significant portion of the training and showcases the extent to which the subspace changes under significant differences in the network parameters and the data distribution. For the sake of completeness and to further highlight the influence of the data distribution on the subspace, we showcase the subspace overlap over the entire duration of the training in LABEL:app:subspace_overlap_for_the_entire_training. As in Section 4.2, we use as subspace dimensionality and refer to Appendix B for the evaluation of different subspace sizes. The analysis results in Figure 3 show that the subspace overlap reduces the further apart the two timesteps and are, but in all cases, the subspace overlap remains significantly above zero, implying that information of previous subspaces can be reused at later timesteps. If the two timesteps are close to each other, the overlap is considerable. Similar to the gradient subspace fraction in Section 4.2, the subspace overlap is often more pronounced for the critic than the actor, particularly for SAC.
@@ 5.2 Potential of gradient subspaces in reinforcement learning
While the network architectures used in reinforcement learning are often small compared to the models used in other fields of machine learning, the dimensionality of the optimization problem is still considerable. Popular optimizers, like Adam (Kingma & Ba, 2014), typically rely only on gradient information, as computing the Hessian at every timestep would be computationally very demanding in high dimensions. However, in Section 4.1, we have seen that the optimization problem is ill-conditioned. Second-order methods, like Newton’s method, are known to be well-suited for ill-conditioned problems (Nocedal & Wright, 1999). With the insights of this paper, it seems feasible to reduce the dimensionality of the optimization problems in RL algorithms by optimizing in the low-dimensional subspace instead of the original parameter space. The low dimensionality of the resulting optimization problems would enable computing and inverting the Hessian matrix efficiently and make second-order optimization methods feasible.
@@ 6 Reproducibility
We applied our analyses to proven and publicly available implementations of the RL algorithms from Stable-Baselines3 (Raffin et al., 2021) on well-known, publicly available benchmark tasks from (Brockman et al., 2016; Plappert et al., 2018b; Tunyasuvunakool et al., 2020). Further experimental details like the learning curves of the algorithms and the fine-grained analysis results for the entire training are displayed in Appendices A and B, respectively. To facilitate reproducing our results, we make our code, as well as the raw analysis data, including hyperparmeter settings and model checkpoints, publically available on the project website.
Appendix A Learning curves
Appendix B Detailed analysis results for all tasks
In Section 4.3, we showed the subspace overlap for a range of timesteps. Appendix B visualizes the subspace overlap criterion with for all future timesteps on tasks that require training for 3 million steps. While in practical applications of gradient subspaces, the subspace would likely be updated multiple times during training, this visualization highlights the influence of the data distribution on the subspace. The plots show a small drop in the subspace overlap for SAC at 1.1 million steps. Since the replay buffer has a size of 1 million samples, this marks the point at which the original data from timestep is completely replaced by new data collected by updated policies. Since the networks’ training data is sampled from the replay buffer, this drop indicates that this change in the data distribution has a negative effect on the subspace overlap. The effect is generally more pronounced for the critic than the actor because the actor’s subspace overlap degrades faster and is already at a relatively low level at the mark of 1.1 million timesteps. For PPO, there is no such drop in the subspace overlap since the algorithm does not use experience replay and instead collects new data for every update.